# A2D: Any-Order, Any-Step Safety Alignment for Diffusion Language Models

### v1

@@ 2 Related Work
As language models become increasingly integrated into real-world applications, ensuring their safe behavior is critical (Hurst et al., 2024; Comanici et al., 2025; Guo et al., 2025). RLHF (Christiano et al., 2017; Ouyang et al., 2022) and DPO (Rafailov et al., 2023) are widely adopted to improve model safety, but they often fail under strong adversarial attacks such as ReNeLLM (Ding et al., 2024) and PAIR (Chao et al., 2025). To enhance robustness, a range of mechanisms have been explored, including external filtering (Lee et al., 2023; Han et al., 2024), representation-level interventions (Zou et al., 2024; Yousefpour et al., 2025), and unlearning-based methods (Barez et al., 2025). Recent studies further show that simple refusal training typically affects only the initial tokens (Qi et al., 2025), highlighting the need for deep alignment across entire generations. In response, new approaches have begun to incorporate safety reasoning beyond surface-level refusals (Jeung et al., 2025b; Zhang et al., 2025a). While these efforts have advanced alignment in autoregressive models, safety alignment for dLLMs remains largely underexplored.
@@ 3 Preliminaries: Diffusion Language Models
In dLLMs, generation proceeds through iterative decoding from a fully masked sequence $\mathbf{x}_{T}=(\texttt{[MASK]},\dots,\texttt{[MASK]})$. At each step $t=T,\dots,1$, the model selects a subset of masked positions; this choice defines the decoding strategy. For each selected index $i$, the model samples
$x_{t-1}^{i}\sim q_{\theta}(\cdot\mid\mathbf{x}_{t}),\quad\text{for all $i$ with }x_{t}^{i}=\texttt{[MASK]},$
@@
Figure 3: Per-token KL divergence between aligned and base dLLMs. Aligned models (LLaDA-Instruct, LLaDA-1.5) vs. Base model (LLaDA-Base) on Harmful BeaverTails under three decoding strategies.
@@ 6.1 Experimental Setup
Table 1: Comprehensive evaluation results on capability and harmfulness for three instruction-tuned dLLMs across four alignment methods. A2D effectively mitigates diverse jailbreak attacks while preserving competitive capability. The top-performing method is shown in bold, and the second-best is underlined.
@@
General
Math
Coding
Zeroshot
PAIR
ReNe
Prefilling
DIJA
Avg.
LLaDA
RT 66.1 37.1 29.9 5.0 66.3 41.9 65.0 80.0 51.6
SFT 65.5 30.8 34.2 43.1 66.7 46.8 57.5 78.1 58.4
VRPO 66.6 40.3 33.5 2.5 31.9 18.8 8.8 46.3 21.7
@@ 6.2 Experimental Results
As shown in Table 1, A2D consistently reduces harmful outputs across all attack types, outperforming all baselines across various dLLMs while preserving model capability. For example, A2D reduces average ASR to 9.0% on LLaDA-1.5 and 2.8% on Dream, whereas all other methods remain above 45% and 20%, respectively. Notably, it brings the ASR of both Prefilling and DIJA close to zero. For DIJA specifically, it achieves 1.3% on LLaDA, 3.8% on LLaDA-1.5, and 0.0% on Dream, demonstrating deep alignment that persists even under partially prefilled completions. A complete breakdown of capability metrics can be found in Table 8.
@@
Figure 4: Per-token KL divergence between A2D-aligned and base dLLMs. Aligned models (LLaDA-1.5, LLaDA-1.5-A2D) vs. Base model (LLaDA-Base) on Harmful BeaverTails under three sampling strategies. LLaDA-1.5-A2D refers to LLaDA-1.5 further aligned with A2D for safety.
@@
PAIR
ReNe
DIJA
PAIR
ReNe
DIJA
PAIR
ReNe
DIJA
PAIR
ReNe
DIJA
PAIR
ReNe
DIJA
@@ 6.4 Additional Analysis on A2D
MMLU
Math
MBPP
PAIR
ReNe
DIJA
@@ A.3 Ablation on Refusal Tokens
A.4 Licenses
@@ C.1 Full Results
Table 8: Full capability benchmark results. Comparison of our method against RT, SFT, and VRPO baselines across General, Math, and Coding benchmarks on LLaDA, LLaDA-1.5, and Dream. Our method consistently matches or outperforms baselines.
@@
RT
SFT
VRPO
Ours
RT
SFT
VRPO
Ours
RT
SFT
VRPO
Ours
  General MMLU 63.3 63.3 64.3 63.2 63.3 63.1 64.1 63.3 70.2 70.2 70.4 70.2
PIQA 74.7 73.9 75.6 74.1 75.0 74.1 75.7 75.7 73.3 73.6 74.2 73.1
Hellaswag 52.8 52.9 53.4 52.9 54.7 52.8 54.7 53.5 53.3 53.3 54.2 53.4
Winogrande 72.4 69.7 71.4 71.7 72.7 69.5 71.5 71.7 71.1 72.5 73.0 71.4
ARC-C 84.3 83.7 85.1 84.2 84.9 84.4 85.8 84.3 60.2 60.2 61.5 59.0
TruthfulQA 49.3 49.7 49.7 50.6 48.6 48.6 50.6 51.0 43.9 41.5 43.9 44.9
@@
  Coding HumanEval 26.8 32.3 29.3 31.7 25.6 30.5 26.2 34.2 52.4 47.6 56.1 56.7
MBPP 33.0 36.0 37.8 38.0 33.8 37.8 38.8 36.4 55.6 54.6 57.2 57.4
@@ C.2 Threshold Analysis
C.3 Qualitative Results

### v2

@@ 2 Related Work
As language models become increasingly integrated into real-world applications, ensuring their safe behavior is critical (Hurst et al., 2024; Comanici et al., 2025; Guo et al., 2025). RLHF (Christiano et al., 2017; Ouyang et al., 2022) and DPO (Rafailov et al., 2023) are widely adopted to improve model safety, but they often fail under strong adversarial attacks such as ReNeLLM (Ding et al., 2024) and PAIR (Chao et al., 2025). To enhance robustness, a range of mechanisms have been explored, including external filtering (Lee et al., 2023; Han et al., 2024), representation-level interventions (Zou et al., 2024; Yousefpour et al., 2025), and unlearning-based methods (Barez et al., 2025). Recent studies further show that simple refusal training typically affects only the initial tokens (Qi et al., 2025), highlighting the need for deep alignment across entire generations. In response, new approaches have begun to incorporate safety reasoning beyond surface-level refusals (Jeung et al., 2025; Zhang et al., 2025a). While these efforts have advanced alignment in autoregressive models, safety alignment for dLLMs remains largely underexplored.
@@ 3 Preliminaries: Diffusion Language Models
In dLLMs, generation proceeds through iterative decoding from a fully masked sequence $\mathbf{x}_{T}=(\texttt{[MASK]},\dots,\texttt{[MASK]})$. At each step $t=T,\dots,1$, the model selects a subset of masked positions; this choice defines the decoding strategy. For each selected index $i$ with $x_{t}^{i}=\texttt{[MASK]}$, the model samples
$x_{t-1}^{i}\sim q_{\theta}(\cdot\mid\mathbf{x}_{t}),$
@@
Figure 3: Per-token KL divergence between aligned and base dLLMs. Aligned models (LLaDA-Instruct, LLaDA-1.5) vs. Base model (LLaDA-Base) on Harmful BeaverTails under three decoding strategies. All results are averaged over 150 harmful prompts from BeaverTails, with shaded regions indicating standard deviation.
@@ 6.1 Experimental Setup
Table 1: Comprehensive evaluation results on capability and harmfulness for three instruction-tuned dLLMs across four alignment methods. A2D effectively mitigates diverse jailbreak attacks while preserving competitive capability. The top-performing method is shown in bold, and the second-best is underlined. All results are averaged over three random seeds, and Original refers to the model without any alignment fine-tuning.
@@
General Math Coding Zeroshot PAIR ReNe Prefilling DIJA Avg.
  LLaDA Original 66.6 41.4 32.6 14.6±0.3 77.5±1.8 56.5±1.3 69.6±2.1 82.9±0.3 60.2
RT 66.2 37.0 29.8 5.0±0.5 65.4±1.2 41.7±0.3 65.8±1.6 79.6±1.0 51.5
SFT 65.3 30.1 34.2 44.0±1.2 69.2±2.4 47.9±0.8 56.9±0.5 78.8±0.9 59.3
VRPO 66.5 40.5 33.5 2.5±0.0 32.3±1.0 19.2±0.6 9.0±1.3 45.0±1.8 21.6
@@ 6.2 Experimental Results
As shown in Table 1, A2D consistently reduces harmful outputs across all attack types, outperforming all baselines across various dLLMs while preserving model capability. For example, A2D reduces average ASR to 9.1% on LLaDA-1.5 and 2.8% on Dream, whereas all other methods remain above 45% and 20%, respectively. Notably, it brings the ASR of both Prefilling and DIJA close to zero. For DIJA specifically, it achieves 1.3% on LLaDA, 3.5% on LLaDA-1.5, and 0.0% on Dream, demonstrating deep alignment that persists even under partially prefilled completions. While there is a slight decrease in general and math performance (often referred to as alignment tax), A2D still outperforms most other methods, remains close to the original model, and delivers substantially stronger safety. A complete breakdown of capability metrics can be found in Table 12.
@@
Figure 4: Per-token KL divergence between A2D-aligned and base dLLMs. Aligned models (LLaDA-1.5, LLaDA-1.5-A2D) vs. Base model (LLaDA-Base) on Harmful BeaverTails under three sampling strategies. LLaDA-1.5-A2D refers to LLaDA-1.5 further aligned with A2D for safety. All results are averaged over 150 harmful prompts from BeaverTails, with shaded regions indicating standard deviation.
@@
PAIR ReNe DIJA PAIR ReNe DIJA PAIR ReNe DIJA PAIR ReNe DIJA PAIR ReNe DIJA
@@ 6.4 Additional Analysis on A2D
MMLU Math MBPP PAIR ReNe DIJA
@@
Table 4: Evaluation results on harmfulness for LLaDA (Nie et al., 2025) across four alignment methods on additional benchmarks. The top-performing method is shown in bold, and the second-best is underlined.
  Method AdvBench ($\downarrow$) JailBreakBench ($\downarrow$)
Zeroshot PAIR ReNe Prefilling DIJA Avg. Zeroshot PAIR ReNe Prefilling DIJA Avg.
  RT 0.0 7.5 57.5 82.5 80.0 45.5 2.5 20.0 57.5 70.0 61.3 42.3
SFT 45.0 76.3 63.8 77.5 82.5 69.0 51.3 87.5 55.0 72.3 63.8 66.0
VRPO 0.0 10.0 13.8 1.3 11.3 7.3 1.3 21.3 16.3 7.5 22.5 13.8
[…]
Table 5: Evaluation results on harmfulness for Dream (Ye et al., 2025) across four alignment methods on additional benchmarks. The top-performing method is shown in bold, and the second-best is underlined.
  Method AdvBench ($\downarrow$) JailBreakBench ($\downarrow$)
Zeroshot PAIR ReNe Prefilling DIJA Avg. Zeroshot PAIR ReNe Prefilling DIJA Avg.
  RT 0.0 13.8 51.3 33.8 81.3 36.0 0.0 16.3 41.3 48.8 67.5 34.8
SFT 16.3 46.3 40.0 70.0 85.0 51.5 27.5 67.5 36.3 68.8 76.3 55.3
VRPO 0.0 10.0 48.8 7.5 60.0 25.3 0.0 7.5 38.8 18.8 45.0 22.0
[…]
@@
Additional evaluations beyond HarmBench.
To further assess robustness of A2D to jailbreaks beyond HarmBench (Mazeika et al., 2024), we evaluate models on two additional benchmarks, AdvBench (Zou et al., 2023) and JailbreakBench (Chao et al., 2024), using both the LLaDA (Nie et al., 2025) and Dream (Ye et al., 2025). The evaluation follows the same experiment setting as Table 1, including three finetuning-based baselines and five representative jailbreak attacks.
As shown in Tables 4 and 5, A2D achieves the lowest average ASR across all jailbreak attacks, benchmarks, and model types, substantially outperforming prior alignment methods. A2D achieves average ASR scores of only 2.8% on AdvBench and 5.3% on JailbreakBench in LLaDA, substantially lower than the next best method, VRPO, which records 7.3% and 13.8% on the same benchmarks. Similarly, on Dream, A2D further reduces the average ASR to below 3.0%, consistently outperforming all baseline methods. These results demonstrate A2D’s robustness across diverse jailbreak settings.
Table 7: Comparison of compliance rates on the long-form benign subset from OR-Bench. The top-performing method is shown in bold.
  Model Original RT VRPO A2D
  LLaDA 55.7 25.7 44.3 47.1
LLaDA-1.5 45.7 28.6 20.0 54.3
Dream 14.3 4.3 12.9 12.9
Robust to Long-Form Benign Prompts.
To assess whether A2D triggers premature refusals in complex benign settings, we evaluate it on a long-form benign subset from OR-Bench (Cui et al., 2025), selecting the 70 longest prompts that appear harmful but are actually safe. As shown in Table 7, A2D attains the highest compliance across models, indicating robust, context-sensitive refusal behavior without false rejections in long-form cases.
@@ 8 Conclusion
Acknowledgements
This work was supported in part by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2024-00457882, AI Research Hub Project), K-CHIPS (Korea Collaborative & High-tech Initiative for Prospective Semiconductor Research) (RS-2024-00405946, 24052-15TC) funded by the Ministry of Trade, Industry & Energy (MOTIE, Korea), and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2025-23525649).
@@ A.3 Ablation on Refusal Tokens
A.4 Computational overhead
Table 8: Training computational cost comparison. A2D matches RT and SFT while being much cheaper than VRPO.
Method RT SFT VRPO A2D
FLOPs (T) 9.7 9.7 29.0 9.7
To estimate the computational cost of A2D, we conduct a single-iteration measurement using the LLaDA-Instruct model fine-tuned with LoRA ($r=32$, $\alpha=64$) applied to the q_proj, k_proj, v_proj, ff_proj, up_proj, and ff_out modules. The setup uses a batch size of 1, input length of 64, and output length of 256 (320 tokens in total).
During training, A2D introduces no additional computational cost compared to standard dLLM fine-tuning. It retains the original masked diffusion objective and only modifies the supervision signal by substituting the [EOS] token for masked harmful spans. As shown in Table 8, A2D requires 9.7T FLOPs, the same as RT and SFT, and substantially lower than VRPO (29.0T FLOPs), which adds extra cost from reinforcement optimization. At inference time, A2D introduces no overhead, following the same decoding process as the base model. It simply monitors the [EOS] signal during generation without requiring additional diffusion steps or classifier guidance.
A.5 Licenses
@@ C.1 Full Results
Table 12: Full capability benchmark results. Comparison of our method against RT, SFT, and VRPO baselines across General, Math, and Coding benchmarks on LLaDA, LLaDA-1.5, and Dream. Our method consistently matches or outperforms baselines. All results are averaged over three random seeds.
@@
RT SFT VRPO Ours RT SFT VRPO Ours RT SFT VRPO Ours
  General MMLU 63.3±0.1 63.2±0.0 64.2±0.0 63.3±0.0 63.3±0.0 63.1±0.0 64.1±0.0 63.3±0.0 70.1±0.0 70.2±0.0 70.3±0.0 70.3±0.1
PIQA 74.5±0.2 74.0±0.3 75.6±0.3 74.3±0.1 75.1±0.1 74.1±0.1 75.6±0.2 75.8±0.1 73.2±0.1 73.7±0.3 74.0±0.2 72.9±0.1
Hellaswag 52.8±0.0 52.8±0.1 53.2±0.2 52.9±0.0 54.8±0.2 52.6±0.1 54.5±0.1 53.6±0.2 53.2±0.1 53.4±0.1 54.2±0.2 53.4±0.2
Winogrande 72.1±0.2 69.4±0.5 71.3±0.2 71.8±0.2 73.1±0.3 69.5±0.0 71.5±0.2 72.1±0.5 70.9±0.4 72.3±0.3 73.4±0.4 71.6±0.4
ARC-C 84.2±0.2 83.3±0.3 84.7±0.5 84.3±0.1 85.2±0.2 84.2±0.2 85.8±0.1 85.2±0.7 60.4±0.3 59.4±0.6 61.5±0.2 59.1±0.2
TruthfulQA 50.3±0.8 49.1±0.5 49.8±0.6 50.5±0.5 49.0±0.3 48.1±0.7 50.7±0.8 51.2±0.4 44.2±0.2 41.4±0.2 44.1±0.5 45.0±0.8
@@
  Coding HumanEval 26.8±0.5 32.3±0.5 29.5±1.3 31.5±0.3 24.8±1.2 29.9±0.9 25.4±0.8 34.8±0.9 52.6±0.3 47.6±0.0 55.7±0.6 57.3±0.9
MBPP 32.8±0.4 36.0±0.2 37.5±0.6 38.4±0.3 33.8±0.1 37.5±0.3 38.8±0.5 36.5±0.1 55.9±0.7 54.5±0.5 57.1±0.1 57.5±0.2
@@ C.2 Threshold Analysis
C.3 Classifier-Based Interpretation of A2D.
A2D can be viewed as implicitly learning a binary classifier $q_{\text{clf}}$ over partially observed contexts. For context $X_{\text{ctx}}$, the training objective encourages the model to approximate
$p_{\theta}(\texttt{[EOS]}\mid X_{\text{ctx}})\approx q_{\text{clf}}(\text{harmful}\mid X_{\text{ctx}}),$
and, for any non-[EOS] token $T$,
$p_{\theta}(T\mid X_{\text{ctx}})\approx q_{\text{clf}}(\text{not harmful}\mid X_{\text{ctx}})\,p_{\theta}^{\text{gen}}(T\mid X_{\text{ctx}}),$
where $p_{\theta}^{\text{gen}}$ denotes the model’s standard generative distribution. This perspective expresses each token probability as a product of a safety factor and a generative factor, with [EOS] acting as a dedicated indicator of harmfulness.
During training, masked positions inside harmful spans are labeled with [EOS], while masked positions in benign spans receive their ground-truth tokens. Because diffusion decoding presents the model with many corrupted partial views of each harmful region, the model repeatedly receives supervision that separates harmful from non-harmful contexts. This repeated conditioning across randomized masks and timesteps yields a stable, distributed decision boundary rather than one that depends on any single token or localized feature.
This interpretation clarifies why [EOS]-based token-level supervision in A2D produces robust and generalizable refusal behavior in diffusion-based language models.
C.4 Qualitative Results
@@
C.5 Adapting A2D to autoregressive LLMs
Table 13: Attack success rates ($\downarrow$) for Qwen-2.5 and LLaMA-3.1 under jailbreak attacks, that a lightweight autoregressive variant of A2D effectively suppresses harmful continuations.
Model ZeroShot PAIR ReNe Prefilling Avg.
Qwen-2.5 2.5 11.9 47.5 47.5 27.4
[…]
LLaMA-3.1 5.0 42.5 41.3 46.3 33.8
[…]
While A2D is motivated by a structural vulnerability unique to diffusion LLMs, namely their any-order, any-step decoding behavior, the underlying mechanism of conditioning on harmful text and enforcing a terminating token is not tied to the diffusion architecture. To examine whether this principle generalizes beyond dLLMs, we conduct a preliminary adaptation of A2D to two autoregressive LLMs, Qwen-2.5-7B (Team, 2024) and LLaMA-3.1-8B (Dubey et al., 2024).
In this AR setting, we adapted A2D to the model by conditioning it on a harmful prefix and training it to predict a special terminating token [EOS] at the subsequent position, with the loss applied only to [EOS] token. For example, when a harmful query appears, the model is trained to emit only [EOS] tokens:
Okay, here is how to make [EOS]
Despite the simplicity of this adaptation, both models exhibit clear reductions in jailbreak attack success rates under strong attacks such as Prefilling, PAIR, and ReNeLLM, as shown in Table 13.
Although a full architectural generalization study is beyond the scope of this work, these preliminary findings indicate that A2D’s core mechanism is not diffusion-specific and may serve as a general framework for suppressing harmful continuations across diverse LLM architectures.
C.6 A2D against [EOS]-interference attack
Table 14: Attack success rates ($\downarrow$) on ReNe variants under A2D alignment. A2D demonstrates strong robustness even under explicit [EOS]-interference attacks.
Model Original A2D-Aligned
ReNe ReNe (+ Long-Answer) ReNe (+ Do-Not-EOS)
LLaDA 56.3 17.5 13.1 12.5
Dream 40.6 9.4 3.8 1.9
To evaluate whether A2D is vulnerable to attacks that directly target its [EOS]-based suppression mechanism, we design two explicit [EOS]-interference attacks. The Long-Answer attack instructs the model to produce the longest possible response, thereby counteracting its tendency to terminate harmful continuations. The Do-Not-EOS attack explicitly tells the model not to output the [EOS] token, directly attempting to override the refusal behavior learned during alignment.
As shown in Table 14, A2D remains robust even under stronger [EOS]-interference attacks. Under the standard ReNe setting, A2D already achieves low attack success rates, and the interference variants yield even lower values. For LLaDA, the interference variants show even lower success rates than the standard ReNe setting: 13.1% under Long-Answer and 12.5% under Do-Not-EOS, compared to 17.5% for ReNe. For Dream, the same pattern holds, with success rates of 3.8% and 1.9% under Long-Answer and Do-Not-EOS, both lower than the ReNe value of 9.4%. These results show that directly targeting [EOS] does not weaken A2D’s refusal behavior; the terminating effect remains stable and often appears more pronounced, indicating that it is not driven by a single [EOS] logit but arises from a distributed suppression pattern across the harmful span.